What will be the remainder when 17^2020is divided by 8

Grade:10

1 Answers

Ritika Das
6 years ago
 
17 is a prime number. Since 17 raised to the power of any number is the multiple of only the factors 17 and 1, we should check its divisibility by 17.
So, 17 divided by 8 gives us a remainder of 1. So the remainder for 172020 divided by 8 will also be 1. You may check this for all numbers like 172, 173, and so on. Divided by 8, they all give remainder 1.
[172=289=(36*8)+1, 173=4913=(614*8)+1]